g(x)=(x+1)² f(x)= 3x-4
find 2g(2) -7f(0)

Answer by stanbon(75887)

g(x)=(x+1)² f(x)= 3x-4
find 2g(2) -7f(0)
-----
2g(2) = 2[2+1]^2 = 2*9 = 18
7f(0) = 7[3*0-4] = 18[-4] = -72
----
2g(2)-7f(0) = 6 - (-72) = 78
==================================
Cheers,
Stan H.
=====================